(a) Expand (a + 1/a)^2
(b) Given that a + 1/a =7, use your result in (a) to find the value of a^2 + 1/a^2

Answer by Alan3354(69443)

(a) Expand (a + 1/a)^2
= a^2 + 2 + 1/a^2
--------------
(b) Given that a + 1/a =7, use your result in (a) to find the value of a^2 + 1/a^2
(a + 1/a)^2 = a^2 + 2 + 1/a^2
7^2 = a^2 + 2 + 1/a^2 = 49
a^2 + 1/a^2 = 47